OpenStudy (anonymous):

Solve. (x+2)(x-10)(x+1)>0. The solution set is...? Or the solution set is all real numbers or there is no solution?

OpenStudy (zarkon):

solve (x+2)(x-10)(x+1)=0 which gives you x=-2,-1,10 pick a number between -infinity,-2 between -2,-1 between -1,10 between 10,infinity plug those numbers into (x+2)(x-10)(x+1) for each one that is positive the entire interval between the roots will also be positive

OpenStudy (anonymous):

\[x \in \mathbb{R}, (-2,1)\cup(10,\infty)\]